OpenStudy (anonymous):

Rewrite problem in equivalent logarithmic form. 1/2=32^-1/5

OpenStudy (campbell_st):

so write 1/2 as 2^{-1} then \[2^{-1} = 32^{-\frac{1}{5}}\] take the log of both sides and remember the log law for powers \[\log(x^a) = a \times \log(x)\] apply this law to the problem